OpenStudy (anonymous):

1. venus revolution 19.4 x 10 ^ -6 s and venus distance to the sun: 108 x 10 ^ 6. so, venus speed of around the sun? 2. planet mass 4 x 10 ^ 20 kg and radius 4,000,000 km. so, how the acceleration of gravity at the surface of the planet? 3. revolution a planet 8 years, the distance of the planet to the sun? 4. 1 ton mass of a car, for 4 second speeds up to 18 m / s. great force that accelerates the car is?

OpenStudy (anonymous):

help me !

OpenStudy (callisto):

Is the speed you need to find in question 1 the angular speed or just distance/time?

OpenStudy (anonymous):

just the speed

OpenStudy (callisto):

Question 1) Distance = \(2\pi r\), r = distance of Venus from the Sun Speed = distance / time

OpenStudy (anonymous):

the other question ?

OpenStudy (callisto):

Question 2) \[F = G\frac{Mm}{r^2}\]\[g= \frac{F}{m} = \frac{ G\frac{Mm}{r^2}}{m} = \frac{GM}{r^2}\]G = gravitational constant.

OpenStudy (anonymous):

:)

OpenStudy (callisto):

Question 3) Same formula as question 1. This time, you need to find r.

OpenStudy (anonymous):

how about number 4?

OpenStudy (callisto):

Question 4) F = ma \[a = \frac{\Delta v}{\Delta t}=\frac{v_1-v_0}{\Delta t}\]

OpenStudy (callisto):

Assume it is frictionless

OpenStudy (anonymous):

@Callisto thank you so much for that :)

OpenStudy (callisto):

Welcome.. Make sure you understand what/why I did there.

OpenStudy (anonymous):

you're very kind @Callisto

OpenStudy (agent0smith):

Did you manage to get 3? 3. revolution a planet 8 years, the distance of the planet to the sun? It's not as simple as @Callisto said. You'll have to equate centripetal force and gravitational force to find r \[\Large \frac{ m v^2 }{ r }= G \frac{ Mm }{ r^2 }\]and also use this to find v to plug into that formula (T is the period of revolution, in seconds) \[\Large T=\frac{ 2 \pi r }{ v }\]